# KKL theorem for the influence of a set of variables

### v1

@@ 6.1. Hypercontractive propositions
[…]
@@ 6.2. Log-Sobolev propositions
where we used an intermediate bound from (23) and the inequality $12^{5d}\beta<$ 1.
@@ 6.3. Coefficients of a $d$-degree function
In this section we prove that the Fourier–Walsh coefficients of a $d$-degree Boolean function lie in the $\mathbb{Z}/2^{d-1}$ lattice. This result was most likely already known, but we could not find reference for it. For results of a similar nature see e.g. Chapter 3 in [9].
@@
A non-trivial example of $f$ satisfying the assumption (26) of small $d$-degree influences is a slightly modified version of the hypertribe function described in Section 5. The nuance to overcome in this case is making the parameter $\alpha$ smaller than $C_{3}$. In order to guarantee this, we may increase the size of the tribes by a constant (depending on $d$). After this modification, the bounds obtained in Proposition 5.3 will be still satisfied for sufficiently large $n$, and following the proof of Proposition 5.5, one can see that $\alpha$ will be suitably small. In this case, for sufficiently large $d$ and $n$, Proposition 5.5 and Proposition 3.6 yield that the approximating function $g$ is the constant $-1$.
@@
Thus, for all $\mathbf{i}\in[n]_{<l}$ we have $\hat{g}(\mathbf{i})=0$, so by Plancharel’s theorem,

### v2

@@ 6.1. Hypercontractive propositions
Let $\mathbf{i}\in[n]_{k+1}$ be arbitrary. Then we have $\operatorname{Inf}_{\mathbf{i}}(g)\leqslant\operatorname{MaxInf}_{k+1}(g)\leavevmode\nobreak\ \leqslant\leavevmode\nobreak\ 1$. Therefore, applying Proposition 4.2 to the function $f=g$, $l=1$ and $\mathbf{i}$ we get
@@ 6.2. Log-Sobolev propositions
where we used an intermediate bound from (20) and the inequality $2^{5d}\beta<1$.
@@ 6.3. Coefficients of a $d$-degree function
In this section we prove that the Fourier–Walsh coefficients of a $d$-degree Boolean function lie in the $\mathbb{Z}/2^{d-1}$ lattice. This result in print has appeared in [9] [Proposition 3.1], but we include our rather short proof to keep the work self-contained.
@@
A non-trivial example of $f$ satisfying the assumption (23) of small $d$-degree influences is a slightly modified version of the hypertribe function described in Section 5. The nuance to overcome in this case is making the parameter $\alpha$ smaller than $C_{3}$. In order to guarantee this, we may increase the size of the tribes by a constant (depending on $d$). After this modification, the bounds obtained in Proposition 5.3 will be still satisfied for sufficiently large $n$, and following the proof of Proposition 5.5, one can see that $\alpha$ will be suitably small. In this case, for sufficiently large $d$ and $n$, Proposition 5.5 and Proposition 3.4 yield that the approximating function $g$ is the constant 1-1.
@@
Thus, for all $\mathbf{i}\in[n]_{<l}$ we have $\hat{g}(\mathbf{i})=0$, so by Plancherel’s theorem,